# Gaussian Process Priors for Boundary Value Problems of Linear Partial Differential Equations

### v1

@@ 5.2 Wave in 3D
Figure 5: Visual comparison between the true solution and B-EPGP prediction for the 3-D wave equation with Neumann boundary conditions on a halfspace. We did not find competing results using Machine Learning techniques for this classic problem. Our older EPGP algorithm ran out of memory on an Nvidia A100. The snapshots are taken in the plane $y=1$.
@@ 5.3 Hybrid B-EPGP in a circular sector
We will also implement our Hybrid B-EPGP method from Section 3.4 for the 2-D wave equation in a circular sector, see Example 3.11. Let $\Omega=\{x,y>0,\,x^{2}+y^{2}\leq 4\}$ and $t\in(0,4)$. The equations are:
$\begin{cases}u_{tt}-(u_{xx}+u_{yy})=0&\text{in }(0,4)\times\Omega\\ u(0,x,y)=f(x,y)&\text{in }\Omega\\ u_{t}(0,x,y)=0&\text{in }\Omega\\ u_{x}(t,0,y)=0&\text{on }(0,4)\times\{0\}\times(0,2)\\ u_{y}(t,x,0)=0&\text{on }(0,4)\times(0,2)\times\{0\}\\ u(t,x,y)=0&\text{on }\text{$x^{2}+y^{2}=4$},\end{cases}$
@@
$e^{at+bx+cy}+e^{at-bx+cy}+e^{at+bx-cy}+e^{at-bx-cy}$
which can be calculated using the ideas of Section G.2. To account for the Dirichlet boundary condition on the circular arc $\Gamma=\{x^{2}+y^{2}=4,\,x,y>0\}$, we simply assign data $u(t_{h},x_{h},y_{h})=0$ at many data points $(t_{h},x_{h}y_{h})\in\Gamma$.
Snapshots of our solution are presented in Figure 7 and the conservation of energy is demonstrated in Figure 6. We iterate that the conservation of energy over time is equivalent to our prediction satisfying both the equation and the boundary condition exactly, see Appendix E.
Figure 6: Proof of energy conservation for the 2-D wave equation in a sector domain. It is clear that our old method S-EPGP leads to dissipation of energy. The Hybrid B-EPGP method performs much better; its error is partly due to our approximation of the energy integral in Appendix E.
@@
Figure 7: Solution of 2-D wave equation in a sector domain calculated using the Hybrid B-EPGP method. The superiority of B-EPGP over EPGP is visually striking and is further evidenced by the energy plot in Figure 6. Animations can be found in the ancillary files sector.mp4 and sectorEPGP.mp4.
5.4 Sampling solutions from data
Our B-EPGP method yields a valid Gaussian process and is hence fully probabilistic. In addition to dealing with noisy data, it can be used to sample solutions of linear PDEs that do not uniquely determined a solution, despite having boundary conditions. This sampling can also be done in the context of given data (noisy or not) of the solution of the PDE. For example, we consider
$\begin{cases}u_{tt}-u_{xx}-u_{yy}=0&\text{for }x,y>0\\ u=0&\text{on }x=0\text{ and }y=0\end{cases}$
and assume given arbitrary data points $(T,X,Y,U)\in\operatorname*{\mathbb{R}}^{D\times 4}$ where $D=25$. These are arbitrary $(T,X,Y)\in[0,4]^{3}$ and $U\in[-10,10]$. We plot snapshots of two random solutions generated using B-EPGP at 5 timepoints in Figure 8, cf. Appendix G.2. This shows the flexibility of our methods when only limited data is available; for comparison, the examples we showed so far are sampled from “enough” data in contexts where the PDE is further constrained to admit unique solution.
(a) $t=0.0$
(b) $t=0.5$
(c) $t=1.0$
(d) $t=1.5$
(e) $t=2.0$
(f) $t=0.0$
(g) $t=0.5$
(h) $t=1.0$
(i) $t=1.5$
(j) $t=2.0$
Figure 8: Irregular/noisy solutions of the 2-D wave equation with Neumann boundary conditions in a quadrant which result from sampling at only 25 data points. The data points are chosen arbitrarily and are the same for both plots. Animations can be found in the ancillary files random1.mp4 and random2.mp4.
Appendix A Calculation of B-EPGP basis for heat and wave equations
A.1 Dirichlet and Neumann conditions on halfspaces
@@
where the boundary conditions are Dirichlet ($B=1$) or Neumann ($B=\partial_{n}$). These are some of the most common boundary conditions used in PDE. As in the body of our paper, $A$ will be assumed to be a single equation and $u$ a scalar field (in Appendix C we will explain the extension to systems and vector fields). By an affine change of variable, +$\mathbb{R}^{n}_{+}$ can be replaced with any other halfspace.
Inspired by the Ehrenpreis-Palamodov Theorem 2.1, we will investigate linear combinations of exponential solutions
@@
which satisfy $A(\partial u)=0$ if
@@
This motivates our definition of the boundary characteristic variety
$V^{\prime}=\{\zeta\in\mathbb{C}^{n-1}\colon\zeta=z^{\prime}\text{ for some }z\in V\},$
as well as giving the ansatz for the B-EPGP basis
@@
A.2 Wave equation
We will calculate the B-EPGP bases for Dirichlet and Neumann for the 2-D wave equation in the domain $y>0$. The calculations extend easily to arbitrary dimensions, but the formulas are cumbersome and we do not include them here. Our calculations extend to all halfspaces parallel to the $t$-axis by affine changes of variable. Considering boundary conditions on halfspaces that are not parallel to the time axis would violate the physical meaning of initial and boundary conditions.
@@ A.2 Wave equation
since for any $(a,b)\in\mathbb{C}^{2}$ there are generically two $c\in\mathbb{C}$ such that $(a,b,c)\in V$.
@@ A.3 Heat equation
since for any $(a,b)\in\mathbb{C}^{2}$ there are generically two $c\in\mathbb{C}$ such that $(a,b,c)\in V$.
@@
which is slightly more general than what we have in Example 3.4 for the 1-D heat equation.
Appendix B Proof that B-EPGP gives all solutions for heat and wave equations
@@
We will only show the calculations in the case of the 2-D wave equation with Dirichlet boundary condition. Increasing the dimension barely changes the argument. The modification required to deal with the heat equation is also minimal (one less integration by parts in time). To deal with the Neumann boundary condition, one uses even extension instead of odd extension in the calculation.
@@
Clearly $\square v(t,x,y)=0$ for $x\neq 0$. We still need to check that $\square v=0$ across $x=0$, but since the odd extension need not have two classical derivatives, we compute the distributional derivative across ${x=0}$. To see this let $\phi\in C_{c}^{\infty}(\{t>0\})$ be a test function. We write $dV=dxdydt$ and integrate by parts
[…]
@@
We next look at the heat and wave equation in rectangles and give a generalization of Theorems 3.8 and 3.10:
Theorem B.1 (Heat and Wave in rectangles).
@@
[…]
@@
Appendix C Ehrenpreis–Palamodov Theorem and EPGP
We will begin with a very precise version of the statement that exponential-polynomial solutions are dense in the space of all solutions for a linear PDE system.
Theorem C.1.
Let $R=\mathbb{C}[\partial]$ and $\Omega\subset\mathbb{R}^{n}$ be an open convex set. Let $A\in R^{\ell\times k}$ and $V=\{z\in\mathbb{C}^{n}\colon\ker A(z)\neq 0\}$. Then there exist a decomposition $V=\bigcup_{i=1}^{m}V_{i}$ into irreducible varieties $V_{i}$ and a set of vector polynomials in $2n$ variables called Noetherian multipliers $\{p_{i,j}\}_{j=1\ldots r_{i},i=1,\ldots m}\subset\mathbb{C}[x,z]^{k}$ such that solutions of the form

### v2

@@ 5.2 High-dimensional Example: 3D Wave Equation
Figure 2: Visual comparison between the true solution and B-EPGP prediction for the 3D wave equation from Section 5.2. The snapshots are taken in the plane $y=1$.
@@ 5.3 Hybrid B-EPGP in a circular sector
We also implement our Hybrid B-EPGP method from Section 3.3 for the 2D wave equation in a circular sector, see Example 3.12. Let $\Omega=\{x,y>0,\,x^{2}+y^{2}\leq 4\}$, $t\in(0,4)$. The equations are:
$\begin{cases}u_{tt}-(u_{xx}+u_{yy})=0&\text{in }(0,4)\times\Omega\\ u(0,x,y)=f(x,y)\text{ and }u_{t}(0,x,y)=0&\text{in }\Omega\\ u_{n}=0&\text{on }xy=0\\ u=0&\text{on }\text{$x^{2}+y^{2}=4$},\end{cases}$
@@
$e^{at+bx+cy}+e^{at-bx+cy}+e^{at+bx-cy}+e^{at-bx-cy}\quad\text{for }a^{2}=b^{2}+c^{2},$
which can be calculated using the ideas of Appendix J.2. To account for the Dirichlet boundary condition on the circular arc $\Gamma=\{x^{2}+y^{2}=4,\,x,y>0\}$, we assign data $u(t_{h},x_{h},y_{h})=0$ at many points $(t_{h},x_{h},y_{h})\in\Gamma$. We give the training details in Appendix M.3.
Snapshots of our solution are presented in Figure 4 and the conservation of energy is demonstrated in Figure 3. We reiterate that the conservation of energy over time is equivalent to our prediction satisfying both the equation and the boundary condition exactly, see Appendix G.
Figure 3: In Section 5.3, EPGP dissipates energy for the 2D wave equation in a sector domain, which is physically incorrect. B-EPGP only has an error due to approximating the energy integral. The same phenomenon is observed for a rectangular domain in Appendix H.1.
@@
Figure 4: Solution of 2D wave equation in a sector domain calculated using the Hybrid B-EPGP method. The superiority of B-EPGP over EPGP is visually striking and is further evidenced by the energy plot in Figure 3. Animations can be found in the supplementary material as sector_B-EPGP.mp4 and sector_EPGP.mp4.
6 Summary
We introduced Boundary Ehrenpreis–Palamodov Gaussian Processes (B-EPGPs), a new framework for constructing GP priors that exactly satisfy systems of linear PDEs with constant coefficients and linear boundary conditions. Our approach generalizes existing methods by leveraging the Ehrenpreis–Palamodov theorem and extending it to include boundary constraints through symbolic algebra and Fourier analysis. Our priors are probability measures on infinite dimensional spaces, the solution sets of our underdetermined PDEs. We provided formal guarantees, explicit constructions, and demonstrated superior empirical performance over state-of-the-art neural operator models and Gaussian process models in both accuracy (by at least a factor of 2) and efficiency (by at least an order of magnitude). Our experiments are implemented in a broad class of domains which includes polygonal and curved boundaries. We do not compare with deterministic numerical solvers as our model is probabilistic.
7 Reproducibility statement
We include our source code and videos of PDE solutions in the supplementary materials.
The comparison to neural operators on the 1D wave equation from Section 5.1 is described in detail in Appendix M.1. The high-dimensional example on 3D wave equation from Section 5.2 is described in detail in Appendix M.2. The hybrid B-EPGP example on the 2D wave equation that has a circular sector domain from Section 5.3 is described in detail in Appendix M.3
The free wave example from Appendix F is described in detail in Appendix M.4. The examples of 2D wave on a different bounded domain from Appendix H and Appendix J are described in detail in Appendix M.5 The comparison between B-EPGP and EPGP on the 2D wave equation from Section I is described in detail in Appendix M.6. The example of 2D heat equation from Appendix K is described in detail in Appendix M.7. The examples of inhomogeneous boundary condition from Appendix L are described in detail in Appendix M.8.
8 Ethics statement
Our method constructs machine learning models constrained to physical equations. While potentially harmful use cases of our models cannot be ruled out, we do not believe that our work currently raises any questions regarding the Code of Ethics.
Appendix A State of the art: machine learning for the 3D wave equation
Many recent related papers either do not tackle wave equations with boundary conditions [36] or do not tackle PDEs in 4D [67]. However, few papers using PINN training and neural operator learning scale to 3D wave physics (time domain and frequency domain), despite the necessity of high computational times. B-EPGP avoids this high computational times.
Neural operators fall into two dominant families. (i) Neural operators (time domain) learn solution operators for the elastic wave PDE from 3D media and source parameters to surface or volumetric wavefields. Architectures include Fourier/Factorized Fourier Neural Operators (FNO/F-FNO) and the U-shaped Neural Operator (UNO). They deliver fast surrogates once trained and generalize across media and sources [52, 53]. (ii) Frequency-domain operator learning replaces the time-domain PDE with the 3D Helmholtz equation; the Helmholtz Neural Operator (HNO) amortizes many forward solves (and even adjoint gradients) across frequencies and media with large speed/memory advantages [89].
For 3D acoustics (Helmholtz), forward PINNs reach FEM-level fields on evaluation grids [73, 72] and report wall-clock comparisons and inference times orders of magnitude below FEM once trained. These studies target low to mid frequencies in room-sized domains, and use DeepXDE/JAX/PyTorch backends with residual and boundary-loss terms.
[…]
Appendix B Gröbner bases
Gröbner bases are a fundamental tool from computational algebra that generalize well-known algorithms in linear algebra and number theory to the setting of multivariate polynomial rings. This section provides an intuitive overview, with a focus on syzygies and how they are used in our construction of boundary-constrained Gaussian processes.
B.1 Gröbner Bases
Gröbner bases extend the idea of the Gaussian elimination algorithm, which solves linear systems, to polynomial systems in several variables. Similarly, they generalize the Euclidean algorithm from univariate polynomials to multivariate settings.
Given a system of polynomial equations with multiple variables, solving or simplifying the system directly is difficult because polynomial rings in several variables are not principal ideal domains. Gröbner bases provide a canonical, algorithmically computable generating set for an ideal of polynomials that simplifies many tasks: solving systems, ideal membership testing, elimination of variables, and finding relations among polynomials.
The key idea is to choose a monomial ordering (such as lexicographic or graded reverse lexicographic order), and then reduce polynomials using a division algorithm adapted to this ordering. A Gröbner basis allows such reductions to behave predictably and mimic row operations in linear algebra.
A fully formal description of Gröbner bases exceeds the scope of this appendix. We refer to the literature for the fully story [78, 20, 1, 29, 26, 13]. Various computer algebra systems implement Gröbner bases, most famously the open source systems Singular [19] and Macaulay2 [28].
B.2 complexity
Sadly, the complexity of Gröbner bases is in the vicinity of EXPSPACE completeness (cf. [60, 61, 6, 62]). Luckily, the “average interesting example” usually terminates instantaneously (less than 0.01 seconds for all examples in this paper), and we could even compute all examples in this paper by hand.
This quick termination is to be expected. Although Gröbner-basis computation has this daunting worst-case bounds, modern algorithms (notably F4/F5) reduce it largely to structured linear algebra on Macaulay-type matrices; on many non-adversarial inputs this step is fast and rarely the overall bottleneck. F4 explicitly organizes reductions as batched Gaussian eliminations, making the cost dominated by dense linear algebra rather than term-by-term reductions [24]. Under generic or semi-regular assumptions, the F5 analysis predicts complexity governed by the degree of regularity, again pointing to linear-algebra–dominated runtimes [4]. In structured tasks such as computing critical points of polynomial maps, generic bounds show that the Gröbner step is tractable (e.g., $D^{O(n)}$), with experiments confirming that it typically behaves well in practice [75]. Worst-case phenomena remain (e.g., dimension-dependent double-exponential behavior), but these arise from carefully engineered instances rather than typical geometric or applied systems [62]. Many geometric and differential-algebraic systems carry symmetries—permutation actions, torus multigradings, or representation-theoretic structure—that dramatically shrink Gröbner computations. For permutation-invariant ideals one can work in the invariant subring or fold equivalent $S$-pairs, collapsing critical pairs and Macaulay matrices [76]. Equivariant Gröbner bases compute a finite basis up to symmetry, solving all finite quotients at once [39]. In toric/sparse settings, Gröbner theory aligns with polyhedral combinatorics; suitable weights give squarefree initial ideals and compact universal bases [77]. Classical GL-invariant families such as determinantal ideals admit symmetry-compatible orders for which the natural generators already form Gröbner bases [12].
In our case, this reduction in complexitiy holds in particular since the Gröbner basis computations only involve the (small) operator equations. They do not have the (potentially big) data as inputs.
B.3 fibres
Let $V\subset\mathbb{C}^{n}$ be given by an ideal $I\subset\mathbb{C}[z^{\prime},z_{n}]$ with $z^{\prime}=(z_{1},\dots,z_{n-1})$, and let $\pi:V\to\mathbb{C}^{n-1}$ be the projection $(z^{\prime},z_{n})\mapsto z^{\prime}$. Following [3], the fibres
$S_{z^{\prime}}\;=\;\{(z^{\prime},z_{n})\in V\}$
can be obtained by a single Gröbner basis computation and then inexpensive specializations. In typical use, once the Gröbner basis is computed, finding the fibre over any concrete $z^{\prime}$ amounts to solving a small zero-dimensional system—fast and routine in computer algebra systems. For implementations of this and a similar algorithm we refer to the command ConstructibleImage in https://homalg-project.github.io/CategoricalTowers/ZariskiFrames/ and to https://github.com/coreysharris/TotalImage.
B.4 Syzygies
In linear algebra, a syzygy is simply a relation among vectors that leads to a zero vector when linearly combined, i.e. an element of the kernel of a matrix with these vectors as columns. These syzygies form a vector space.
In the more general context of polynomial systems, syzygies are relations among (vectors of) polynomials that annihilate a given combination. More precisely, given a tuple of (vectors of) polynomials $f_{1},\ldots,f_{m}$, a syzygy is a tuple of polynomials $(g_{1},\ldots,g_{m})$ such that $\sum_{i}g_{i}f_{i}=0$. The set of all such relations forms a module over the polynomial ring, called the syzygy module.
Computing a basis of this module is analogous to solving a homogeneous system of linear equations, but over a multivariate polynomial ring [^2] This is very explicitly written in the formulas (7) and (8).. Instead of using the Gaussian algorithm to compute a basis, one uses Buchberger’s algorithm to compute a Gröbner basis. Then, one can (in a non-trivial way called Schreyer’s algorithm) read of the syzygies.
B.5 Syzygies for B-EPGP
In our construction of boundary-constrained Ehrenpreis–Palamodov Gaussian processes (B-EPGP), we seek linear combinations of exponential-polynomial functions that satisfy both the differential equation and the boundary conditions exactly. These basis functions are of the form $e^{z\cdot x}$, where $z$ lies in the characteristic variety associated with the PDE system.
The boundary condition introduces additional linear constraints on these exponentials. Specifically, for a fixed fiber of frequencies $S_{z^{\prime}}$ (i.e., a set of vectors $z$ sharing the same spatial components), we need to find coefficients $w_{z}$ such that a linear combination $\sum_{z\in S_{z^{\prime}}}w_{z}B(z)$ vanishes, where $B(z)$ represents the boundary operator evaluated at $z$.
This is precisely a syzygy computation: we seek all tuples $(w_{z})_{z\in S_{z^{\prime}}}$ such that a symbolic linear combination of the boundary terms gives zero. These syzygies are computed by standard Gröbner basis algorithms, and the resulting linear combinations form the building blocks for our B-EPGP basis functions.
We comment on the role of the frequency vectors $z$ in this computation. For each fixed $S_{z^{\prime}}$, the values $z\in S_{z^{\prime}}$ are concrete complex numbers and the $B(z)$ are concrete complex vectors. Hence, the $w_{z}$ for these concrete numbers might be computed by a Gaussian algorithm. However, this induces massively redundant computations of the Gaussian algorithm for each frequency vector in each step of the training algorithm, including backpropagation through this computation. Hence, it is advantageous to symbolically precompute the $w_{z}$ for symbolic $z$.
For a correct computation, we need to include the polynomial relations between the $z\in S_{z^{\prime}}$, i.e. that they are distinct fibers of $z^{\prime}$ of the map $V\to\mathbb{C}^{n-1}$. The Gröbner basis needs to be computed over the residue class ring of the polynomial ring by these polynomial relations.
Gröbner bases and syzygies enable a symbolic guarantee that all samples from our Gaussian process model satisfy the PDE system and the boundary conditions exactly, not just approximately. This exactness is crucial in applications such as physics-informed modeling or solving ill-posed PDEs, where small violations of boundary constraints can lead to qualitatively incorrect behavior.
Appendix C Calculation of B-EPGP basis for heat and wave equations
C.1 Dirichlet and Neumann conditions on halfspaces
@@
where the boundary conditions are Dirichlet ($B=1$) or Neumann ($B=\partial_{n}$). These are some of the most common boundary conditions used in PDE. As in the body of our paper, $A$ will be assumed to be a single equation and $u$ a scalar field (in Appendix E we will explain the extension to systems and vector fields). By an affine change of variable, $\mathbb{R}^{n}_{+}$ can be replaced with any other halfspace.
Inspired by the Ehrenpreis–Palamodov Theorem 2.1, we will investigate linear combinations of exponential solutions
@@
which satisfy $A(\partial)u=0$ if
@@
This motivates our definition of the boundary characteristic constructible set
$V^{\prime}=\{\zeta\in\mathbb{C}^{n-1}\colon\zeta=z^{\prime}\text{ for some }z\in V\}.$
Generically, this constructible set is a variety, which we then call boundary characteristic variety. This gives the ansatz for the B-EPGP basis
@@
C.2 Wave equation
We will calculate the B-EPGP bases for Dirichlet and Neumann for the 2D wave equation in the domain $y>0$. The calculations extend easily to arbitrary dimensions, but the formulas are cumbersome and we do not include them here. Our calculations extend to all halfspaces parallel to the $t$-axis by affine changes of variable. Considering boundary conditions on halfspaces that are not parallel to the time axis, which would violate the physical meaning of initial and boundary conditions.
@@ C.2 Wave equation
since for any $(a,b)\in\mathbb{C}^{2}$ there is at least one (and generically two) $c\in\mathbb{C}$ such that $(a,b,c)\in V$.
@@ C.3 Heat equation
since for any $(a,b)\in\mathbb{C}^{2}$ there is at least one (and generically two) $c\in\mathbb{C}$ such that $(a,b,c)\in V$.
@@
which is slightly more general than what we have in Example 3.5 for the 1D heat equation.
Appendix D Proof that B-EPGP gives all solutions for heat and wave equations
@@
We will only show the calculations in the case of the 2D wave equation with Dirichlet boundary condition. Increasing the dimension barely changes the argument. The modification required to deal with the heat equation is also minimal (one less integration by parts in time). To deal with the Neumann boundary condition, one uses even extension instead of odd extension in the calculation.
@@
Clearly $\square v(t,x,y)=0$ for $x\neq 0$. We still need to check that $\square v=0$ across $x=0$, but since the odd extension need not have two classical derivatives, we compute the distributional derivative across x=0{x=0}. To see this let $\phi\in C_{c}^{\infty}(\{t>0\})$ be a test function. We write $dV=dxdydt$ and integrate by parts
$\int v\square\phi dV$ $=\int_{x>0}u\square\phi dV-\int_{x<0}u(t,-x,y)\square\phi(t,x,y)dV$
$=-\int_{x=0}u\phi_{x}dydt-\int_{x>0}u_{t}\phi_{t}-u_{x}\phi_{x}-u_{y}\phi_{y}dV-\int_{x=0}u\phi_{x}dydt$
$+\int_{x<0}u_{t}(t,-x,y)\phi_{t}(t,x,y)+u_{x}(t,-x,y)\phi_{x}(t,x,y)-u_{y}(t,-x,y)\phi_{y}(t,x,y)dV$
$=-\int_{x=0}u_{x}\phi dydt+\int_{x>0}\phi\square udV+\int_{x=0}u_{x}\phi dydt+\int_{x<0}\phi\square udV=0,$
@@
We next look at the heat and wave equation in rectangles and give a generalization of Theorems 3.9 and 3.11:
Theorem D.1 (Heat and Wave in rectangles).
@@
Only the last assertion is non-obvious and requires a careful distributional calculation as in the proof of Theorem 3.6 above, taking $\phi\in C_{c}^{\infty}(\mathbb{R}^{1+2})$ and showing that $\int u\square\phi dV=0$ by careful integration by parts. We omit the details.
@@
Appendix E Ehrenpreis–Palamodov Theorem and EPGP
We will begin with a very precise version of the statement that exponential-polynomial solutions are dense in the space of all solutions for a linear PDE system. For notational clarity, we write $\mathbb{C}[\partial]:=\mathbb{C}[\partial_{1},\ldots,\partial_{n}]$, $x=(x_{1},\ldots,x_{n})$, and $z=(z_{1},\ldots,z_{n})$.
Theorem E.1.
Let $\Omega\subset\mathbb{R}^{n}$ be an open convex set. Let $A=A(\partial)\in\mathbb{C}[\partial]^{\ell\times k}$ be an operator matrix of a system of linear PDEs with constant coefficients and $V=\{z\in\mathbb{C}^{n}\colon\ker A(z)\neq 0\}$ its characteristic variety. Then there exist a decomposition $V=\bigcup_{i=1}^{m}V_{i}$ into irreducible varieties $V_{i}$ and a set of vector polynomials in $2n$ variables called Noetherian multipliers $\{p_{i,j}\}_{j=1\ldots r_{i},i=1,\ldots m}\subset\mathbb{C}[x,z]^{k}$ such that solutions of the form